OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through the point (4, -1) and is perpendicular to the line whose equation is 2x - y - 7 = 0?

OpenStudy (anonymous):

y = -1/2x + 1 y = 1/2x - 1 y = 2x + 1

OpenStudy (lgbasallote):

it is perpendicular to the line y = 2x - 7 so the slope of the perpendicular line is -1/2 now use point slope formula y - (-1) = -1/2 (x-4) y + 1 = -1/2 (x-4) y + 1 = -1/2 x + 2 y = -1/2 x + 2 - 1 y = -1/2 x + 1

OpenStudy (anonymous):

thanks